Tags:
create new tag
view all tags

Shift 20101116

17:00 AA, LC, LD, LN

Plan

  • Load optics from set10nov11_1414.dat set file.
  • Check detector attenuation.
  • Tune for low background.
  • Wirescans to determine trajectory.
  • Iterate until trajectory and background OK.
  • Aim to get beam at LW IP below 20Ám.
  • Timing (hopefully using Naito-san's super scope).
  • Vertical alignment using manipulator and knowledge of laser position.
  • Horizontal alignment using manipulator notch.
  • Vertical scans iterated over various horizontal positions.
  • Retiming if necessary.
  • Achieve collisions.
  • Celebrate in bar.

Start

A lot of changes to the damping ring were made and the previous double ATF2 shift had considerable trouble getting a decent beam in the extraction line.

AA tuning linac / beam transport for transmission into the damping ring

LD tuning extraction line

  • Loaded optics from set10nov11_1414.dat set file.
  • Background large due to generally poor beam quality.

LD notes here.

Wirescans / Background

OTR screens

The chambers for the OTR measurements (OTR0, OTR1 upstream of LWIP, OTR2, OTR3 downstream) sit very low and might cause background for our optics. The chambers have stay positions ("5V") well out of the beam and can be accessed through the front end OTR program. When the SLAC team left the chambers were in their OTR measurement low positions (e.g. OTR0 Y = -283 um) but can be set out further from the beam and the SLAC team set them out for us remotely. Readings:

  • OTR0 Y = 2723 Ám
  • OTR1 Y = 7271 Ám
  • OTR2 Y = 2019 Ám
  • OTR3 Y = 3883 Ám

Wirescans / Background

LC wire scan details here along with skew quad scan

20:00 Background around 4000 with 14dB of attenuation.

Wire Scanner Dimension Centre (mm) Width (Ám)
MW1X X 19.403▒0.167 433.0▒230.3
Y 53.089▒0.023 62.7▒25.2

QK2X I (A) MW1X - Yσ
0.00 62.7▒25.2
-4.5 57.5▒23.0
-6.5 52.9▒22.6
-8.5 53.8▒18.3

Set back to -6.5A as that had the smallest Y σ.

Wire Scanner Dimension Centre (mm) Width (Ám)
MW2X X NA NA
Y 61.987▒0.037 95.2▒39.5

Magnet polarity change and aggressive optics

9.20pm Beam off and the KEK technicians went to change the polarity of QM16FF. This is required for the set of optics designed for us by Mark Woodley, to get a predicted vertical size of ~ 2Ám at the LWIP.

9.45pm Magnet changed - takes about 30 mins so needs to be swapped back at 12.30am for the next shift.

10pm LD loaded MW's optics (10nov162202.dat) - beam on. Initially all beam lost in extraction line, LW detector saturated. LD put more attenuation in, up to 31dB. AA tuning beam.

After steering, attenuation reduced again to 14dB, background ~ 11000 counts. Tuning continued.

Background tuned down to ~ 6000 - 7000 counts. Wire scan to check these new optics

Wire Scanner Dimension Centre (mm) Width (Ám)
MW1X X 20.422▒0.088 446.0▒104.2
Y 52.870▒0.009 26.1▒10.5
MW2X X 34.477▒0.005 68.4▒5.2
Y 62.269▒0.005 26.2▒5.9

Tuned to move the horizontal focus nearer the IP.

Wire Scanner Dimension Centre (mm) Width (Ám)
MW1X X 20.327▒0.088 399.1▒101.3
Y 52.773▒0.007 25.2▒7.4
MW2X X 34.512▒0.009 104.0▒10.8
Y 61.868▒0.004 22.6▒4.7

These tuned MW optics are saved in 10nov17_0005.dat

Timing

00.35 LD, LN optimise timing.

Collision Search

00.40 Horizontal scan with screen set to laser position - at ~ -170 chamber position horizontal. Vertical scan between 1200 and 1250 chamber position.

Scanning between 900 and 1450 vertically - no signal.

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r6 - 18 Nov 2010 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback